# Deep RL Needs Deep Behavior Analysis: Exploring Implicit Planning by Model-Free Agents in Open-Ended Environments

### v1

@@ B.2 Behavioral Phase Segmentation via Unsupervised Clustering
To understand how agents modulate their movement patterns across time and task demands, we applied unsupervised clustering to the agent’s locomotion features. We used the bayesmove package to identify three latent movement states based on turning angle and step size over a 7-timestep moving window [81]. We next used conditional inference trees to identify which task variables best predicted transitions between these states [82].
@@
Figure 7: Agents exhibit distinct movement patterns aligned with task variables such as food, fatigue, and predator presence. (A) Variable importance from conditional inference trees [82] predicting transitions between three movement states identified via unsupervised clustering [81]: short-, mid-, and long-range locomotion. Key predictors include food level, positional uncertainty, predator presence, and nearby cow density. (B) A representative episode segment colored by inferred movement state. (C) The same trajectory overlaid with predator encounters and eat events. State 1 (mid-range) frequently coincides with predator presence, while state 3 (short-range) clusters near food and eat actions. These results suggest that DRL agents flexibly shift between movement regimes in response to internal drives and environmental context—without requiring weight updates.
@@ B.3 Comparison to Real Insect Search Patterns
Figure 8: Agent exploration trajectories qualitatively resemble real-world search patterns in ants and bees. (A) Adapted from [68]: an ant’s search trajectory following nest displacement, showing outwardly expanding loops over time. (B) Adapted from [83]: bee trajectories during their first, second, and third departures from the nest. In both cases, the animals exhibit looping paths that rotate azimuthally and expand in spatial scale—closely matching the trajectories observed in DRL agents during early foraging. These parallels suggest that similar search strategies may emerge across biological and artificial systems when operating under uncertainty and partial observability.
@@ B.4 GLM Outputs and Multicollinearity Checks
B.5 Training Dynamics and Cross-Entropy Evolution
To better understand how behavior evolves over training, we tracked the average policy entropy and the cross-entropy between agent actions and greedy choice baselines. Figure 9 shows how decision confidence increases over time, as reflected in the entropy and cross-entropy dynamics. Early in training, agents exhibit high entropy and low cross-entropy, consistent with undirected exploration. As learning progresses, cross-entropy rises and entropy declines, indicating increasing policy stability and confidence in selecting goal-directed actions.
@@ B.5 Training Dynamics and Cross-Entropy Evolution
Figure 9: Cross-entropy and entropy dynamics reveal a transition from undirected exploration to confident, goal-directed behavior. Evolution of average cross-entropy and entropy over training for PPO-RNN agents. Early in training, agents display low cross-entropy (vs. greedy choice) and high entropy, consistent with exploratory behavior. As training progresses, entropy declines while cross-entropy increases, reflecting increased confidence and policy refinement.
@@ B.6 Memory and Predator-Related Ablations
Figure 11: Disabling the auxiliary position-prediction objective removes structured spatial representations. Training dynamics and movement patterns for PPO-RNN agents without the auxiliary path integration objective. The absence of structured spatial uncertainty plots (top left) reflects the removal of the predicted position output normally produced by the auxiliary loss.
@@
Figure 13: GLM results confirm that patch revisitation decisions are shaped by uncertainty, reward history, and predator presence. (A) Full GLM output underlying Figure 3, obtained using statsmodels.formula.api in Python. The model classified patches as chosen (1) or not (0) based on average historical variables, using agent ID as a fixed effect. A single model was fit jointly across five PPO-RNN agents, spanning 7978 revisitation decisions. (B) Variance inflation factor (VIF) scores show no excessive multicollinearity among predictors (all VIF $<$ 10).
Figure 14: Front-facing agents exhibit denser local exploration in early training. Each row shows a representative early-exploration trajectory for an agent with a front-facing field of view (FOV), as opposed to the 360 egocentric view used in the baseline configuration. Front-facing agents tended to perform more localized and dense exploratory behavior during the early phases of training.
Figure 15: Front-facing FOV agents exhibit earlier onset of long-range exploration. Variant of Figure 4 for PPO-RNN agents with a front-facing field of view. While data is available only for early and mid training, these agents showed a faster shift toward long-range exploration compared to baseline agents. This modification may serve as a useful manipulation in future studies of sensory constraints.
@@ C.3 Regularization and Train-Test Split for Decoding
Appendix D Toward Architectures for Cognitive Map Formation (Future Directions)
@@
While our agents do not yet use structured maps, our findings suggest that such architectures may enhance both behavior and interpretability. Building on these findings about memory, planning, and spatial encoding in model-free agents, we highlight one architectural direction and outline broader opportunities for future work. An important next step is to explore architectures that more closely reflect mammalian spatial navigation capabilities. In biological systems, both allocentric and egocentric representations support planning and memory [84], particularly through grid-based spatial computations [85].
@@
These “cognitive maps” [5, 86, 87, 88, 89] extend beyond current position encoding—they support retrospection, prospective planning, and flexible decision-making [90, 91, 92]. Recent models have begun incorporating cognitive map-like representations into reinforcement learning frameworks [93].
@@
In animals, spatial maps are modulated by task variables such as goals, landmarks, and reward-related events [94, 95, 96, 97]. These advantages have been recognized in machine learning as well, with grid-based representations increasingly integrated into transformers [98] and actor-critic agents [99] to improve task performance, efficiency, and interpretability.
@@
Appendix E Craftax licensing permissions

### v2

@@ B.2 Behavioral Phase Segmentation via Unsupervised Clustering
To understand how agents modulate their movement patterns across time and task demands, we applied unsupervised clustering to the agent’s locomotion features. We used the bayesmove package to identify three latent movement states based on turning angle and step size over a 7-timestep moving window [cullen2022identifying]. We next used conditional inference trees to identify which task variables best predicted transitions between these states [hothorn2015ctree].
@@
Figure 7: Agents exhibit distinct movement patterns aligned with task variables such as food, fatigue, and predator presence. (A) Variable importance from conditional inference trees [hothorn2015ctree] predicting transitions between three movement states identified via unsupervised clustering [cullen2022identifying]: short-, mid-, and long-range locomotion. Key predictors include hunger level (food), positional uncertainty (uncertainty), predator presence (enemy_present), and nearby cow density (num_passives_nearby). (B) A representative episode segment colored by inferred movement state. (C) The same trajectory overlaid with predator encounters and eat events. State 1 (mid-range) frequently coincides with predator presence, while state 3 (short-range) clusters near food and eat actions. These results suggest that DRL agents flexibly shift between movement regimes in response to internal drives and environmental context—without requiring weight updates as these results were explored on test arenas with fixed weights.
@@ B.3 Comparison to Real Insect Search Patterns
Figure 8: Agent exploration trajectories qualitatively resemble real-world search patterns in ants and bees. (A) Adapted from [wehner1981searching]: an ant’s search trajectory following nest displacement, showing outwardly expanding loops over time. (B) Adapted from [osborne2013ontogeny]: bee trajectories during their first, second, and third departures from the nest. In both cases, the animals exhibit looping paths that rotate azimuthally and expand in spatial scale—closely matching the trajectories observed in DRL agents during early foraging. These parallels suggest that similar search strategies may emerge across biological and artificial systems when operating under uncertainty and partial observability.
@@ B.4 GLM Outputs and Multicollinearity Checks
B.5 Training Dynamics and Entropy Evolution
To better understand how behavior evolves over training, we tracked the average policy entropy, Figure 9 shows how decision confidence increases over time, as reflected in the entropy dynamics that indicate increasing policy stability and confidence in selecting goal-directed actions.
@@ B.5 Training Dynamics and Entropy Evolution
Figure 9: Entropy dynamics reveal a transition from undirected exploration to confident, goal-directed behavior. Evolution of average entropy over training for PPO-RNN agents. As training progresses, entropy declines and the log probability of the selected actions increases, reflecting increased confidence and policy refinement. Entropy slowly increases later in training as the agent discovers ways to maximize both task performance and policy entropy, reflective of the loss function.
@@ B.6 Memory and Predator-Related Ablations
Figure 11: Disabling the auxiliary position-prediction objective limits the predator killing abilities of agents. Training dynamics and movement patterns for PPO-RNN agents without the auxiliary path integration objective. The absence of structured spatial uncertainty plots (top left) is due to the removal of the predicted position output normally produced by the auxiliary loss, since predicted current position is not outputted by the network in this experiment. The primary difference in learning history dynamics relative to the baseline model is that without the position-prediction objective, the emergence of predator killing abilities of agents is worse. This ability requires agents crafting a tool, with a sword being the highest damage tool, and reducing the predator health to 0. Predator killing is difficult because the predators can also kill the agent.
@@
Figure 13: GLM results confirm that patch revisitation decisions are shaped by uncertainty, reward history, and predator presence. (A) Full GLM output underlying Figure 3, obtained using statsmodels.formula.api in Python. The model classified patches as chosen (1) or not (0) based on average historical variables, using agent ID as a fixed effect. A single model was fit jointly across five PPO-RNN agents, spanning 7978 revisitation decisions. (B) Variance inflation factor (VIF) scores show no excessive multicollinearity among predictors (all VIF << 10).
Figure 14: Front-facing agents exhibit denser local exploration in early training. Each row shows a representative early-exploration trajectory for an agent with a front-facing field of view (FOV), as opposed to the 360 egocentric view used in the baseline configuration. Front-facing agents tended to perform intermittent localized and dense exploratory behavior during the early phases of exploration.
Figure 15: Front-facing FOV agents exhibit earlier onset of long-range exploration. Variant of Figure 4 for PPO-RNN agents with a front-facing field of view. These agents showed a faster shift toward long-range exploration compared to baseline agents. This modification may serve as a useful manipulation in future studies of sensory constraints.
@@ C.3 Regularization and Train-Test Split for Decoding
Appendix D Generalized linear model for predicting neural activity from current position
As a supplementary analysis, we explored how position was represented at the single neuron level in the RNNs. NeMoS (Neural ModelS) was used for fitting generalized linear models (GLMs) that predicted neural activity from the current position of the agent, see https://github.com/flatironinstitute/nemos for extended documentation [nemos2025]. One GLM was fit per neuron, but multiple episodes (arena configurations) were used for each model fit, typically 5-10 episodes minimum to avoid overtraining to one arena configuration. Prior to setting up the model, the 96 x 96 arena was coarse-grained into 14 x 14 position bins, and a binary variable per bin was encoded that reported whether or not the agent was in a given bin (196 total regressor coefficients). Neural data was also shifted into the positive axis (lowest value zeroed), and normalized to the range of 0-1 per neuron to extract the activity changes relative to a given neuron’s baseline for easier cross-neuron comparison.
The GLM fit the first 70% of each episode in the training set of episodes Figure 19. The majority of neurons fit had comparable performance between the test and training sets Figure 20.
Figure 19: Example test/train split for a single episode to be used in GLM fitting. GLM models, one model per neuron, with 5-10 episodes per model, predicted neural activity from position. The first 70% of an episode was trained, while the last 30% was the test split.
Figure 20: Log likelihood scores for the same GLM model (single neuron fit), showing the difference in log likelihood between test and train data, one model per data point. The data here is a PPO-RNN baseline run.
Approximately 100/512 neurons had good fit performance using just position regressors across runs, with values typically in the range of 60-120 position-sensitive neurons per training run. An example good fit is shown in Figure 21.
We then plotted the number of neurons with GLM coefficients in the three largest mean-valued k-means clusters out of five clusters, see Figure 22 for the selection. We also plotted the average GLM coefficient per position bin Figure 23. The number of strongly position encoding neurons increased with radial distance from the origin of the arena, and the average GLM coefficient magnitude also increased with radial distance from the origin Figure 24. This trend was seen in all main conditions, PPO-RNN with and without path integration, and PQN-RNN. Thus, an accumulation circuit may ramp with distance from origin and pressure the agent to return to the origin the farther away it gets, partially explaining the periodic origin revisiting behavior observed across conditions. This pattern was present even without the position predicting auxiliary objective Figure 25 and Figure 26, and also in the PQN-RNN runs Figure 27 and Figure 28, suggesting it is fundamental to the task solution. Example scripts to use this analysis method are in the github repository for this paper.
Figure 21: Example of a single neuron, multi-episode model fit. Each of the 512 GLM models fit, one per neuron, has one GLM coefficient for each of 196 position bins (14 x 14) that the arena grids were coarse-grain binned into for this analysis. We show an example good fit for a single neuron’s data (black), fit to 12 different episodes for one model (fit in red).
Figure 22: Number of position-sensitive neurons per position bin in the arena for PPO-RNN. Each of the 512 GLM models fit, one per neuron, has one GLM coefficient for each of 196 position bins (14 x 14) that the arena grids were coarse-grain binned into for this analysis. The heatmap of the number of well-fit neurons per position bin is shown in the plot. The data here is a PPO-RNN baseline run. The 0’s in the uppermost row and rightmost column of bins are due to the agents never going out to those position bins in these runs.
Figure 23: Heatmap of average GLM coefficient magnitude per position bin for PPO-RNN. Each of the 512 GLM models fit, one per neuron, has one GLM coefficient for each of 196 position bins (14 x 14) that the arena grids were coarse-grain binned into for this analysis. These plots show the average GLM coefficient value across well-fit neurons per bin (left), and the Gaussian smoothed shape of the same heatmap (right). The data here is a PPO-RNN baseline run. The 0’s in the lowermost row and rightmost column of bins are due to the agents never going out to those position bins in these runs.
Figure 24: Average position bin GLM coefficient versus radial distance from the arena origin (center) for PPO-RNN. Each of the 512 GLM models fit, one per neuron, has one GLM coefficient for each of 196 position bins (14 x 14) that the arena grids were coarse-grain binned into for this analysis. The average GLM coefficient for the well-fit neurons per radial bin distance from the origin was plotted, finding the neural response to a position change increases with distance from the origin. The data here is a PPO-RNN baseline run. Data points are the mean across neurons while error bars are 95% CI.
Figure 25: Heatmap of average GLM coefficient magnitude per position bin for PPO-RNN without path integration. Each of the 512 GLM models fit, one per neuron, has one GLM coefficient for each of 196 position bins (14 x 14) that the arena grids were coarse-grain binned into for this analysis. These plots show the average GLM coefficient value across well-fit neurons per bin (left), and the Gaussian smoothed shape of the same heatmap (right). The data here is a PPO-RNN without position-prediction (no path integration) run. The 0’s in the lowermost row and rightmost column of bins are due to the agents never going out to those position bins in these runs.
Figure 26: Average position bin GLM coefficient versus radial distance from the arena origin (center) for PPO-RNN without path integration. Each of the 512 GLM models fit, one per neuron, has one GLM coefficient for each of 196 position bins (14 x 14) that the arena grids were coarse-grain binned into for this analysis. The average GLM coefficient for the well-fit neurons per radial bin distance from the origin was plotted, finding the neural response to a position change increases with distance from the origin. The data here is a PPO-RNN without position-prediction (no path integration) run. Data points are the mean across neurons while error bars are 95% CI.
Figure 27: Heatmap of average GLM coefficient magnitude per position bin for PQN-RNN. Each of the 512 GLM models fit, one per neuron, has one GLM coefficient for each of 196 position bins (14 x 14) that the arena grids were coarse-grain binned into for this analysis. These plots show the average GLM coefficient value across well-fit neurons per bin (left), and the Gaussian smoothed shape of the same heatmap (right). The data here is a PQN-RNN run. The 0’s in the lowermost row and rightmost column of bins are due to the agents never going out to those position bins in these runs.
Figure 28: Average position bin GLM coefficient versus radial distance from the arena origin (center) for PQN-RNN. Each of the 512 GLM models fit, one per neuron, has one GLM coefficient for each of 196 position bins (14 x 14) that the arena grids were coarse-grain binned into for this analysis. The average GLM coefficient for the well-fit neurons per radial bin distance from the origin was plotted, finding the neural response to a position change increases with distance from the origin. The data here is a PQN-RNN run. Data points are the mean across neurons while error bars are 95% CI.
Appendix E Alternative DRL Objective: Parallelized Q-networks (PQN)
Recent work has shown that the off-policy RL algorithm parallelized Q-networks (PQN) can perform well in vanilla Craftax [gallici2024simplifying]. We explored whether this alternative DRL algorithm performs differently compared to PPO on ForageWorld. We adapted the PQN algorithm to ForageWorld using the LSTM model architecture and default hyperparameters from [gallici2024simplifying], which differ in some ways from our PPO implementation (for example, using an LSTM instead of a GRU for the RNN). Except where noted in Table 3, we used the same environment and training hyperparameters between PPO and PQN. This experiment provides an orthogonal test of ForageWorld– would a different algorithm, model, and training parameters result in different behavior?
We found that the PQN-LSTM models performed comparably to the PPO-GRU baseline for the subset of PQN models that learned the task (Figure 29), but that only a subset of PQN runs would converge to this level of performance– across the two successful test conditions that varied the exploration hyperparameters of PQN (in particular controlling the rate of decay of $\epsilon$, the chance of taking a random action instead of the optimal action under the current Q value function of PQN), only  30% of PQN runs performed comparably to the PPO model, with the remainder failing catastrophically in a way that was not seen in the baseline PPO runs. We believe this is the result of suboptimal exploration by PQN, i.e. the poorly performing PQN runs learned simple and reliable but suboptimal behaviors early in training and then were slow to deviate from them later on. For the runs that did perform comparably to PPO, we found overall similar learning histories and pathing behavior to the PPO baseline runs (Figure 30, Figure 31), but with the noticeable behavioral differences of much lower rates of predator killing (i.e. reducing predator HP to 0) (Figure 31) in the learning history analysis. Thus, this result suggests that, despite showing comparable overall performance at the task, PPO seems to converge to a predator fighting strategy, while PQN converges to a predator evasion strategy.
Figure 29: Performance comparison for PQN versus PPO. Individual PQN-LSTM runs performed comparably to the PPO-GRU baseline for two different exploration hyperparameter configurations, but the majority of runs failed to learn more than a trivial policy. Condition 1 used the $\epsilon$ decay parameters shown in Table 3, while condition 2 adjusted $\epsilon$ start, finish, and decay parameters so that $\epsilon$ decayed half as fast and ended at a final value of 0.1 (twice the chance of a random action). No condition resulted in uniformly successful runs. Other exploration hyperparameter values were tested, which are detailed in Table 3, but did not result in any high-performing runs.
Figure 30: Representative trajectories from three episodes for PQN-RNN. Each row (A–C) illustrates three sequential stages from a single episode: early exploratory loops (left), a path sampled during the transition to revisitation (middle), and the full episode trajectory with revisited patch regions highlighted in orange (right).
Figure 31: Training dynamics and movement patterns learning history for PQN-RNN agents with the auxiliary path integration objective. PQN had overall similar patterns to PPO, but was found to not learn predator killing as well (gray arrow).
Appendix F Toward Architectures for Cognitive Map Formation (Future Directions)
@@
While our agents do not yet use structured maps, our findings suggest that such architectures may enhance both behavior and interpretability. Building on these findings about memory, planning, and spatial encoding in model-free agents, we highlight one architectural direction and outline broader opportunities for future work. An important next step is to explore architectures that more closely reflect mammalian spatial navigation capabilities. In biological systems, both allocentric and egocentric representations support planning and memory [vorhees2014assessing], particularly through grid-based spatial computations [moser2014grid].
@@
These “cognitive maps” [epstein2017cognitive, behrens2018cognitive, todd2020foraging, ambrogioni2023rethinking, piray2024reconciling] extend beyond current position encoding—they support retrospection, prospective planning, and flexible decision-making [csicsvari2007place, olafsdottir2018role, denovellis2021hippocampal]. Recent models have begun incorporating cognitive map-like representations into reinforcement learning frameworks [piray2021linear].
@@
In animals, spatial maps are modulated by task variables such as goals, landmarks, and reward-related events [hardcastle2015environmental, hardcastle2017multiplexed, boccara2019entorhinal, sosa2023hippocampal]. These advantages have been recognized in machine learning as well, with grid-based representations increasingly integrated into transformers [whittington2021relating] and actor-critic agents [banino2018vector] to improve task performance, efficiency, and interpretability.
@@
Appendix G Craftax licensing permissions
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We tried to carefully ensure the strength of evidence matched the strength of claims in our work.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have discussed limitations in section 5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not have theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Environment, model and analysis codes are provided in the paper’s github repository (https://github.com/RileySE/Craftax-Foraging/tree/foraging).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes] .
  24. Justification: Environment, model and analysis codes are provided in the paper’s github repository (https://github.com/RileySE/Craftax-Foraging/tree/foraging). We provide key hyperparameters and other settings to reproduce experiments in Table 3
  25. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  26. 6.
    Experiment statistical significance
  27. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  28. Answer: [Yes]
  29. Justification: Yes, when needed to justify statistical claims, the authors use 95% CI as error bars unless otherwise specified.
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  30. 7.
    Experiments compute resources
  31. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  32. Answer: [Yes] .
  33. Justification: Yes, see the appendix.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  35. 8.
    Code of ethics
  36. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  37. Answer: [Yes]
  38. Justification: The authors have reviewed the code of ethics and believe our work is not in violation of any rules.
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  39. 9.
    Broader impacts
  40. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  41. Answer: [Yes]
  42. Justification: Yes, see section 5.
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  43. 10.
    Safeguards
  44. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  45. Answer: [N/A] .
  46. Justification: We did not scrape any data or release a model at significant risk of dual-use applications.
  47. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  48. 11.
    Licenses for existing assets
  49. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  50. Answer: [Yes] .
  51. Justification: Our work modifies the Craftax environment, whose free use license is given in the appendix: Appendix G. Additionally, the proper license terms were followed from the respective host journals for the adapted, previously published figures in Figure 8. We paid for the use of figure [wehner1981searching], and the other figure was published under an unrestricted use license, Creative Commons Attribution License, as long as the author is accredited [osborne2013ontogeny].
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  52. 12.
    New assets
  53. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  54. Answer: [Yes] .
  55. Justification: We have released the ForageWorld environment code and accompanying model and analysis code, which are provided on Github at (https://github.com/RileySE/Craftax-Foraging/tree/foraging).
  56. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  57. 13.
    Crowdsourcing and research with human subjects
  58. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  59. Answer: [N/A] .
  60. Justification: We do not perform human research.
  61. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  62. 14.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  63. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  64. Answer: [N/A] .
  65. Justification: We do not perform human research.
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  66. 15.
    Declaration of LLM usage
  67. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  68. Answer: [N/A] .
  69. Justification: We do not use LLMs.
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.